Mrs cotter recently bought a new smart car and wants to test its street racing capabilities. She starts her car and slams on the gas while driving across the staff parking lot after school. If her car starts at rest and she accelerated at a rate of 2.34 m/s, then how long would it take her in seconds to reach a speed pfb26.82 m/s (60 mph)


1
Expert's answer
2020-09-15T13:54:58-0400

a=∆v/∆t

∆t=∆v/a = 26.82-2.34/2.34 = 10,4615385 s.